# 1]FAIR at Meta 2]Univ. Grenoble Alpes, Inria, CNRS, Grenoble INP, LJK, France

### v2

@@ 3.1 OneFlow Scales Better than AR
Table 1: Parity FLOP ratio in a controlled experiment. Both models were trained on 2B image-text pairs. Parity FLOP Ratio represents the relative amount of OneFlow FLOPs needed to match the final AR + FM performance.
@@ 3.2 Mixed Modal Pre-training Enables Better Generation and Understanding
In this section, we study the impact of mixed-modal pretraining by comparing concurrent and sequential training approaches. Specifically, we train two 1B variants: one using sequential pretraining (either text-to-image or image-to-text), and another using the same data but with 20% of examples incorporating concurrent generation.
Our ablation experiment reveals that mixed-modal training significantly improves downstream visual question answering performance (Figure 6) and provides a minor boost to image generation. These results demonstrate the benefits of concurrent mixed-modal pretraining. Given these consistent improvements, we use the mixed-modal pretrained model for all subsequent finetuning experiments.
@@ 3.5.1 Classifier-free guidance improves text detailedness
CFG Scale 0.0 1.0 1.4 2.0 AR GPT4o Score 5.36 5.61 5.53 5.44 5.60
Figure 10: Edit Flows with classifier-free guidance produces longer and more detailed answers, improving metrics that involve VLM as a judge (Cheng et al., 2025). Additional examples of classifier-free guidance effects on text generation are shown in Figure 16.
@@
As shown in Figures 10 and 16, higher CFG values consistently increase the length and detail of generated text. We quantitatively evaluated caption quality and hallucination using CapArena (Cheng et al., 2025) by prompting GPT4-o. Our findings show that increasing CFG leads to more detailed captions, with OneFlow matching AR’s level of detail at a guidance scale of 1. However, this increased detail comes at the expense of hallucinations at very high CFG values.

### v3

@@ 3.1 OneFlow Scales Better than AR
Table 1: Parity FLOPS ratio in a controlled experiment. Both models were trained on 2B image-text pairs. Parity FLOPS Ratio represents the relative amount of OneFlow FLOPS needed to match the final AR + FM performance.
@@ 3.2 Mixed Modal Pre-training Enables Better Generation and Understanding
In this section, we study the impact of mixed-modal pretraining by comparing concurrent and sequential training approaches. Specifically, we train two 1B variants: one using sequential pretraining (either text-to-image or image-to-text), and another using the same data but with 20% of examples incorporating concurrent generation. Our ablation experiment reveals that mixed-modal training significantly improves downstream visual question answering performance (Figure 6) and provides a minor boost to image generation. These results demonstrate the benefits of concurrent mixed-modal pretraining. Given these consistent improvements, we use the mixed-modal pretrained model for all subsequent finetuning experiments.
3.2.1 Mixed Generation Probability
Figure 8: Effect of mixed-generation probability on performance. Higher mixed-generation probabilities consistently improve image understanding (CIDEr, ROUGE-L) and compositional generation (DPG Score), while image generation quality (FID) remains stable. This demonstrates that mixed training enhances understanding and compositional capabilities without compromising generation performance.
In this section, we study the impact of different mixed-generation probabilities over 200k training steps. Results show a clear monotonic trend: higher mixed-generation probability consistently improves image understanding across all metrics. CIDEr scores increase from 135.3 (0% mixed) to 138.3 (100% mixed), ROUGE-L improves from 0.67 to 0.71, and DPG Score rises from 0.67 to 0.71. Importantly, image generation quality (FID, DPG Score) remains stable across all probabilities, demonstrating that mixed training strengthens understanding capabilities without degrading generation performance.
@@ 3.5.1 Classifier-free guidance improves text detailedness
Figure 11: OneFlow with classifier-free guidance produces longer and more detailed answers OneFlow matches the AR baseline in GPT-4o as judge (Cheng et al., 2025), while higher guidance scales trade off between detailedness and hallucination. CHAIRS and CHAIRI measure the rate of hallucinated objects at the sentence and instance levels, respectively (lower is better).
@@
Figure 12: Edit Flows with classifier-free guidance produces longer and more detailed answers, examples of classifier-free guidance effects on text generation are shown in Figure 18.
@@
As shown in Figures 12 and 18, higher CFG values consistently increase the length and detail of generated text. We quantitatively evaluated caption quality and hallucination using CapArena (Cheng et al., 2025) by prompting GPT4-o. Figure 11 shows that increasing CFG leads to more detailed captions, with OneFlow matching AR’s level of detail at a guidance scale of 1. However, this increased detail comes at the expense of hallucinations at very high CFG values.
@@ B.4 Sampling Steps on Caption Performance
B.5 t-Independent Parameterization
Time Location CIDEr
none 1.28
begin_of_seq 1.27
begin_of_text 0.33
Table 6: CIDEr scores for different time embedding locations.
We empirically found t-independent parameterization to perform better despite theoretical motivation for t-dependence. This likely occurs because $X_{t}$ already encodes sufficient information about the noise level. We ablated the location of time token insertion: begin_of_seq places it at the beginning of the sequence, begin_of_text at the start of text tokens, and none corresponds to t-independent parameterization.
B.6 K-scheduler
Figure 22: Effect of $k$-scheduler degree. Using the schedule $\kappa_{t}=t^{k}$, we find that linear scheduling ($k=1.0$) achieves the best results. Higher degrees lead to overly aggressive token deletion.
We conducted ablations over different k-schedulers. As described in (Gat et al., 2024), there are three variants of the k-scheduler: linear, quadratic, and cubic. We find that the linear scheduler works best.